OpenStudy (anonymous):

Using a sheet of graph paper, solve the following system of equations graphically. Be sure to show any work, use a straight edge, and be neat. When you are finished, enter the solution set in the box below and turn in your graph paper to your teacher. x + y = 0 x - y + 2 = 0 I do not understand what to do?

OpenStudy (solomonzelman):

\(\Large\color{red}{ \bf x + y = 0 }\) \(\Large\color{red}{ \bf x - y = -2 }\)

OpenStudy (solomonzelman):

If it doesn't seem logical to just see the answer... you can add the equations, solve for x, and then find the y.

OpenStudy (solomonzelman):

Sorry, graphically-:( graph each equation, and the point where they intersect is the solution.

OpenStudy (anonymous):

i do not understand how to graph it when it is in the terms x+y=0 and x-y=-2 @SolomonZelman

OpenStudy (solomonzelman):

\(\Large\color{red}{ \bf x+y=0 ~~~~~~~~~~ x-y=-2 }\) \(\Huge\color{blue}\downarrow\) \(\Huge\color{blue}\downarrow\) \(\Large\color{red}{ \bf y=-x ~~~~~~~~~~~~ -y=-x-2 }\) \(\Huge\color{blue}\downarrow\) \(\Large\color{red}{ \bf~~~~~~~~~~ y=x+2 }\)

OpenStudy (solomonzelman):

graph each equation, it will not be hard, promise :)

OpenStudy (anonymous):

okay thanks!!! @SolomonZelman